Two blocks of mass m=25 kg and M= 20 kg are connected via pulley with a configuration. Assuming that the surface are smooth and the angle is 45 degrees. Find the acceleration blocks and tension of the cable.
1
Expert's answer
2020-10-05T10:55:45-0400
"Ma=Mg-T\\ma=T-mg\sin{45}\\(m+M)a=g(M-m\sin{45})\\(25+20)a=9.8(20-25\sin{45})\\a=0.51\frac{m}{s^2}"

Tension is

"20(0.51)=20(9.8)-T\\T=186\ N"